# Rademacher’s Theorem for Calderon-Zygmund-type Spaces

### v1

@@ Abstract
Rademacher’s Theorem is a classical result stating that a Lipschitz function on $\mathbb{R}^{d}$ possesses a total differential almost everywhere. This implies that if $f$ is a function defined on $\mathbb{R}^{d}$ belonging to the Calderón–Zygmund space $T^{\infty}_{1}(x)$ for every $x\in\mathbb{R}^{d}$, then $f\in t^{\infty}_{1}(x)$ for almost every $x\in\mathbb{R}^{d}$.
The main purpose of this paper is to extend this property to a broader functional framework. More precisely, we replace the Lipschitz condition by the assumption that $f$ belongs to a general weighted Calderón–Zygmund space $T^{p}_{\phi}(x)$ for $x\in E$, where $E$ is a measurable subset of $\mathbb{R}^{d}$ and $\phi$ is a weight with fractional indices. We then show that, under suitable assumptions, the function belongs to $t^{p}_{\phi}(x)$ for almost every $x\in E$.
The result can be obtained either by imposing additional hypotheses or by working within a suitably adapted $t^{p}_{\phi}$ space. Whenever relevant, we also provide counterexamples illustrating the sharpness and the limitations of the statement.
@@ 1 Introduction
The $T^{p}_{u}$ spaces were initially introduced by Calderón and Zygmund [3] : given a point $x_{0}\in\mathbb{R}^{d}$, $p\in[1,\infty]$ and a number $u\geq-d/p$, $T^{p}_{u}(x_{0})$ denotes the class of functions $f$ in $L^{p}(\mathbb{R}^{d})$ for which there exists a polynomial $P$ of degree strictly less than $u$ with the property that
@@
$r^{-d/p}\|f-P\|_{L^{p}(B(x_{0},r))}=o(r^{u})\text{ when }r\rightarrow 0^{+},$ (2)
then $f\in t^{p}_{u}(x_{0})$. They allow to determine the pointwise regularity of functions that are not locally bounded (see for example [6, 16, 9]). Indeed, since the spaces $T^{p}_{u}$ are decreasing with $u$, one may define the $p$-exponent of $f$ :
@@
If $p=\infty$, $T^{p}_{u}(x_{0})$ is the classical pointwise Hölder space, denoted by $\Lambda^{u}(x_{0})$.
Rademacher’s Theorem is a classical result which states that a Lipschitz function on $\mathbb{R}^{d}$ has a total differential almost everywhere. This means that if $f$ is a function defined on $\mathbb{R}^{d}$ which belongs to the space $T^{\infty}_{1}(x)$ for every $x\in\mathbb{R}^{d}$, then $f\in t^{p}_{1}(x)$ for almost every $x\in\mathbb{R}^{d}$, where the polynomial involved in the definition of the spaces $t^{p}_{1}(x)$ has degree less than or equal to 1.
A. Calderón generalized the differentiability of Sobolev functions by showing that if $\Omega\subset\mathbb{R}^{d}$ is bounded and open, then every function in $W^{p}_{1}(\Omega)$ is differentiable almost everywhere for $p>d$ [5]. This result follows naturally from the Lebesgue differentiation theorem and the Sobolev embedding theorem, and Rademacher’s theorem appears as a special case, since every Lipschitz function on $\Omega$ lies in $W^{\infty}_{1}(\Omega)$. Rademacher’s Theorem also admits a broader formulation for Lipschitz mappings from Euclidean spaces into arbitrary metric spaces, where differentiability is defined in terms of metric differentials rather than classical derivatives [4].
The original idea here was to replace the Lipschitz condition in Rademacher’s Theorem by the assumption that the function belongs to $T^{p}_{\phi}(x)$ for $x\in E$, where $E$ is a measurable subset of $\mathbb{R}^{d}$, and to show that the function then belongs to $t^{p}_{\phi}(x)$ for almost all $x\in E$, where space $T^{p}_{\phi}$ (resp. $t^{p}_{\phi}$) is a generalized version of $T^{p}_{u}$ (resp. $t^{p}_{u}$) by replacing $r\mapsto r^{u}$ appearing in (1) (resp. (2)) by a function parameter (see section 2). This yields an $L^{p}$ version of Rademacher’s Theorem, and this has been established for the spaces $T^{p}_{u}$, with $u\in\mathbb{N}_{0}$ [17].
Here, $\mathcal{L}$ stands for the Lebesgue measure, $\mathbb{N}_{0}=\{0,1,2,3,$...\} and $\mathbb{N}=\{1,2,3,..\}$.

### v2

@@ Abstract
Rademacher’s Theorem can be interpreted as an almost-everywhere little-$o$ improvement principle: if a function admits a uniform pointwise first-order Lipschitz control at every point, then this control improves to a vanishing one at almost every point. In the language of Calderón–Zygmund pointwise spaces, this means that
$f\in T^{\infty}_{1}(x)\quad\forall x\in\mathbb{R}^{d}\qquad\Longrightarrow\qquad f\in t^{\infty}_{1}(x)\quad\text{for a.e. }x\in\mathbb{R}^{d}.$
The purpose of this paper is to establish an analogous almost-everywhere improvement principle in a refined $L^{p}$ setting. We consider pointwise Calderón-Zygmund spaces $T^{p}_{\phi}(x)$ defined via polynomial approximation in $L^{p}$ with a function parameter $\phi$, allowing for fractional regularity indices and logarithmic corrections through Boyd functions. We prove that, under natural assumptions on $\phi$, the uniform membership
$f\in T^{p}_{\phi}(x)\quad\forall x\in E$
on a measurable set $E\subset\mathbb{R}^{d}$ implies an almost-everywhere improvement to a vanishing approximation rate, namely
$f\in t^{p}_{\phi,n+1}(x)\quad\text{for a.e. }x\in E,$
where $n<\underline{b}(\phi)\leq\overline{b}(\phi)<n+1$.
The proof combines measurability arguments, a generalized Whitney extension theorem, and fine properties of Sobolev spaces. We also show that this result is essentially sharp: in general, one cannot expect almost-everywhere membership in $t^{p}_{\phi,n}(x)$ for fractional indices, and explicit counterexamples are provided.
@@ 1 Introduction
Pointwise Calderón–Zygmund spaces provide a flexible framework to quantify local regularity through polynomial approximation in $L^{p}$. They were originally introduced by Calderón and Zygmund [3]. Given a point $x_{0}\in\mathbb{R}^{d}$, $p\in[1,\infty]$ and a real number $u\geq-d/p$, the space $T^{p}_{u}(x_{0})$ consists of functions $f\in L^{p}(\mathbb{R}^{d})$ for which there exists a polynomial $P$ of degree strictly less than $u$ such that
@@
$r^{-d/p}\|f-P\|_{L^{p}(B(x_{0},r))}=o(r^{u})\quad\text{when }r\rightarrow 0^{+},$ (2)
then $f$ is said to belong to the space $t^{p}_{u}(x_{0})$.
These spaces allow one to describe the pointwise regularity of functions that may fail to be locally bounded (see for instance [6, 16, 9]). Since the spaces $T^{p}_{u}(x_{0})$ are decreasing with respect to $u$, one can define the $p$-exponent of a function $f$ at $x_{0}$ by
@@
When $p=\infty$, the space $T^{\infty}_{u}(x_{0})$ coincides with the classical pointwise Hölder space $\Lambda^{u}(x_{0})$.
From this point of view, Rademacher’s Theorem can be interpreted as an almost-everywhere improvement principle. It states that if a function is uniformly Lipschitz, then its first-order approximation improves from a bounded rate to a vanishing one at almost every point. In terms of Calderón–Zygmund spaces, this means that
$f\in T^{\infty}_{1}(x)\quad\forall x\in\mathbb{R}^{d}\qquad\Longrightarrow\qquad f\in t^{\infty}_{1}(x)\quad\text{for a.e. }x\in\mathbb{R}^{d},$
where the polynomial involved in the definition of $t^{\infty}_{1}(x)$ has degree at most $1$.
This principle has been extended in several directions. In particular, Calderón showed that if $\Omega\subset\mathbb{R}^{d}$ is bounded and open, then every function in the Sobolev space $W^{p}_{1}(\Omega)$ is differentiable almost everywhere whenever $p>d$ [5]. This result follows from the Lebesgue differentiation theorem together with Sobolev embeddings, and Rademacher’s Theorem appears as a limiting case, since Lipschitz functions belong to $W^{\infty}_{1}(\Omega)$. Rademacher-type results also exist in more general metric settings, where differentiability is formulated in terms of metric differentials [4].
The purpose of this paper is to establish an analogous almost-everywhere improvement principle in a refined $L^{p}$ framework. More precisely, we replace the Lipschitz assumption by a pointwise control in generalized Calderón–Zygmund spaces $T^{p}_{\phi}(x)$, where the power function $r^{u}$ in (1) is replaced by a function parameter $\phi$. These spaces allow for fractional regularity indices and finer scales involving logarithmic corrections. They are defined using Boyd functions and will be introduced in detail in Section 2.
Given a measurable set $E\subset\mathbb{R}^{d}$, the guiding question of this work is the following:
If a function $f$ satisfies $f\in T^{p}_{\phi}(x)$ for every $x\in E$, does this uniform pointwise control improve to a vanishing one at almost every point of $E$?
For integer indices, this question was answered positively by Ziemer in [17] for the spaces $T^{p}_{u}(x)$ with $u\in\mathbb{N}_{0}$, yielding an $L^{p}$ version of Rademacher’s Theorem.
Our strategy follows the classical approach of Ziemer, adapted to the function-parameter setting. First, we show that the map
$x\longmapsto\|f\|_{T^{p}_{\phi}(x)}$
is measurable. Using Lusin’s Theorem, this allows us to localize the problem on compact subsets where uniform bounds are available. We then rely on a generalized Whitney extension theorem to construct a smooth extension agreeing almost everywhere with $f$ on these subsets. Fine properties of Sobolev functions are subsequently used to obtain an almost-everywhere vanishing approximation rate for the extension, which is finally transferred back to the original function.
For fractional indices, this approach yields an essentially optimal result. In general, one can only expect almost-everywhere membership in the space $t^{p}_{\phi,n+1}(x)$, where
$n<\underline{b}(\phi)\leq\overline{b}(\phi)<n+1,$
and where the polynomial appearing in the definition of the space has degree at most $n+1$. We further provide explicit counterexamples showing that this threshold cannot, in general, be lowered.
Throughout the paper, $\mathcal{L}$ denotes the Lebesgue measure, $\mathbb{N}_{0}=\{0,1,2,3,\dots\}$ and $\mathbb{N}=\{1,2,3,\dots\}$.